OpenStudy (anonymous):

Is the following an arithmetic sequence? 5, 0, -5, -10,... Yes or No

hartnn (hartnn):

if there is a common difference between the terms, then it is.

hartnn (hartnn):

is 0-5 = -5-0 = -10-(-5) ??

OpenStudy (♪chibiterasu):

Common difference meaning that when subtracting each one, they result in the same difference.

OpenStudy (anonymous):

ok so what do I subtract it by?

OpenStudy (♪chibiterasu):

Well, we go backwards.

OpenStudy (♪chibiterasu):

-10 -(-5) -5 - 0 0 - 5 Do they all have the same difference?

OpenStudy (♪chibiterasu):

Well, let's see. -10 -(-5) -(- always cancels out to and addition so: -10 + 5 = ? -10 + 5 = -5 Now, if this is an arithmetic sequence we have to prove that the rest of the difference are equal to -5.

OpenStudy (♪chibiterasu):

So, for it to be an arithmetic sequence: -5 - 0 must be proven to be -5 0 - 5 must be proven to be -5 -5 - 0 = ? -5 - 0 = -5 0 - 5 = ? 0 - 5 = -5 -5 -5 -5 Now can we say this is an arithmetic sequence?

OpenStudy (anonymous):

Yes

OpenStudy (♪chibiterasu):

There you go :3